References

[Adams] J.F. Adams, Infinite Loop Spaces, Annals Studies in Math., vol. 90, Princeton University Press, Princeton, 1978.

[And72] D. W. Anderson, Relationship among K-theories, Lecture Notes in Math., vol. 341, Springer-Verlag, 1973, pp. 57–72.

[AT65] S. Araki and H. Toda, Multiplicative structures in mod q cohomology theories I, Osaka J. Math. 2 (1965), 71–115.

[Art] V. Artamonov, Projective Modules over Group Rings of Nilpotent Groups, Algebra, Moskov. Gos. Univ. (1982), 7–23 (Russian); English transl. in AMS Translations, vol. 154, 1992, pp. 11–24.

[Atiy56] M. Atiyah, On the Krull–Schmidt theorem with application to sheaves, Bull. Soc. Math. France 84 (1956), 307–317.

[Atiy57] M. Atiyah, Vector bundles over an elliptic curve, Proc. London Math. Soc. 7 (1957), 414–452.

[Atiyah] M. Atiyah, K-theory, Benjamin, New York, 1967.

[AM] M. Atiyah and I. McDonald, Introduction to Commutative Algebra, Addison-Wesley, 1969.

[AT] M. Atiyah and D. Tall, Group Representations, λ-rings and the J-homomorphism, Topology 8 (1969), 253–297.

[BA] N. Jacobson, Basic Algebra II, Freeman & Co., 1980.

[Bak] A. Bak, K-theory of Forms, Annals Studies in Math., vol. 98, Princeton University Press, Princeton, 1981.

[Ban] G. Banaszak, Algebraic K-theory of number fields and rings of integers and the Stickelberger ideal, Annals of Math. 135 (1992), 325–360.

[BG] G. Banaszak and W. Gajda, Euler systems for higher K-theory of number fields, J. Number Theory 58 (1996), 213–252.

[BP71] M. Barratt and S. Priddy, On the homology of non-connected monoids and their associated groups, Proc. Symp. Pure Math., vol. 22, 1971, pp. 31–36.

[Bass] H. Bass, Algebraic K-theory, Benjamin, 1968.

[Bass76] H. Bass, Euler characteristics and characters of discrete groups, Inv. Math. 35 (1976), 155–196.

[B72] H. Bass, Unitary algebraic K-theory, Lecture Notes in Math., vol. 342, Springer-Verlag, 1973.

[BHS] H. Bass, A. Heller and R. Swan, The Whitehead group of a polynomial extension, Publ. Math. IHES 22 (1964), 61–79.

[BMS] H. Bass, J. Milnor and J.-P. Serre, Solution of the congruence subgroup problem for SL_{n} (n ≥ 3) and Sp_{2n} (n ≥ 2), Publ. Math. IHES 33 (1967), 59–137.

[BM] H. Bass and M.P. Murthy, Grothendieck groups and Picard groups of abelian group rings, Annals Math 86 (1967), 16–73.

[BT] H. Bass and J. Tate, The Milnor ring of a global field, Lecture Notes in Math., vol. 342, Springer-Verlag, 1973, pp. 349–446.

[BBD] A. Beilinson, J. Berstein, and P. Deligne, Faisceaux pervers, Asterisque vol. 100,, 1982.

[Berrick] J. Berrick, An Approach to Algebraic K-theory, Pitman, 1982.

[Bl74] S. Bloch, K_{2} and algebraic cycles, Annals of Math. 99 (1974), 349–379.

[Bl86] S. Bloch, Algebraic cycles and higher K-theory, Adv. in Math. 61 (1986), 267–304.

[BK] S. Bloch and K. Kato, p-adic etale cohomology, Publ. IHES 63 (1986), 107–152.

[BL] S. Bloch and S. Lichtenbaum, A spectral sequence for motivic cohomology, unpublished 1995 preprint, available at www.math.uiuc.edu/K-theory/0062.

[Bor] A. Borel, Cohomologie réelle stable de groupes S-arithmétiques classiques, C. R. Acad. Sci. (Paris) Sér. A-B (1972), A1700–A1702; Cohomology of Arithmetic Groups, Proc. 1974 ICM Vancouver, 1975, pp. 435–442.

[Bor1] A. Borel, Stable real cohomology of arithmetic groups, Ann. scient. Éc. Norm. Sup. 7 (1974), 235–272.

Typeset by AMS-TEX [BoSe] A. Borel and J.-P. Serre, Le Théorème de Riemann-Roch (d’après Grothendieck), Bull.

Soc. Math. de France 86 (1958), 97–136.

[BSh] Z. Borevich and I. Shafarevich, Number Theory, Benjamin, New York, 1967. [B-AC] N. Bourbaki, Algèbre commutative, Chap.1-Chap.9, Hermann, Paris, 1961-83.

[BF] A. K. Bousfield and E. M. Friedlander, Homotopy theory of Γ-spaces, spectra, and bisimplicial sets, Lecture Notes in Math., vol. 658, Springer-Verlag, 1978, pp. 80–130.

[Br] W. Browder, Algebraic K-theory with coefficients ℤ/p, Lecture Notes in Math., vol. 657, Springer-Verlag, 1978.

[BrwS] J. Browkin and A. Schinzel, On Sylow 2-subgroups of K_{2}(O_{F}) for quadratic number fields F, J. Reine Angew. Math. 331 (1982), 104–113.

[BG] K. Brown and S. Gersten, Algebraic K-theory as generalized sheaf cohomology, Lecture Notes in Math., vol. 341, Springer-Verlag, 1973, pp. 266–292.

[Bu12] J. Buhler, R. Crandall, R. Ernvall, T. Metsänkylä and M. Shokrollahi, Irregular primes and cyclotomic invariants to 12 million, J. Symbolic Comput. 31 (2001), 89–96.

[BH] J. Buhler and D. Harvey, Irregular primes to 163 million, preprint, arXiv:0912.2121 (2008).

[Carter] D. Carter, Lower K-theory of finite groups, Comm. in Alg. 8 (1980), 1927–1937.

[CCMT] J. Caruso, F. Cohen, J. P. May and L. Taylor, James maps, Segal maps, and the Kahn-Priddy theorem, Trans. Amer. Math. Soc. 281 (1984), 243–283.

[CLL] K. Choo, T. Lam and E. Luft, On free products of rings and the coherence property, Lecture Notes in Math., vol. 342, Springer-Verlag, 1973, pp. 135–143.

[Co] H. Cohen, A course in computational algebraic number theory, Graduate Texts in Mathematics 138, Springer-Verlag, 1993.

[Cohen] M. Cohen, A Course in Simple–Homotopy Theory, Springer-Verlag, 1973.

[Cohn65] P.M. Cohn, Universal Algebra, Harper&Row, 1965.

[Cohn66] P.M. Cohn, Some remarks on the invariant basis property, Topology 5 (1966), 215–228.

[CHSW] G. Cortiñas, C. Haesemeyer, M. Schlichting and C. Weibel, Cyclic homology, cdhcohomology and negative K-theory, Annals of Math. 167 (2008), 549-563.

[CHWW] G. Cortiñas, C. Haesemeyer, M. Walker and C. Weibel, Bass’ NK groups and cdhfibrant Hochschild homology, Inventiones Math. 181 (2010), 421-448.

[DS] R.K. Dennis and M. Stein, K_{2} of radical ideals and semilocal rings revisited, Lecture Notes in Math., vol. 342, Springer-Verlag, 1973.

[Drin] V. Drinfeld, Infinite-dimensional vector bundles in algebraic geometry: an introduction, Progress in Math., vol. 244, Birkhäuser, Boston, 2006, pp. 263–304.

[Dror] E. Dror, Acyclic spaces, Topology 11 (1972), 339–348.

[DS] J. Dupont and C.-H. Sah, Scissors Congruences II, J. Pure Applied Algebra 25 (1982), 159–195.

[DM1] W. Dwyer and S. Mitchell, On the K-theory spectrum of a smooth curve over a finite field, Topology 36 (1997), 899–929.

[EGA] A. Grothendieck and J. Dieudonné, Eléments de Géométrie Algébrique, Publ. Math. IHES; I 4 (1960); II 8 (1961); III 11 (1961); 17 (1963); IV 20; (1964); 24 (1965); 28 (1966); 32 (1967); I (revised) (1971), Springer-Verlag.

[EGS] P. Elbaz-Vincent, H. Gangl and C. Soulé, Quelques calculs de la cohomologie de GL_{N} (ℤ) et de la K-théorie de ℤ, C. R. Acad. Sci. (Paris) 335 (2002), 321–324.

[FJ] T. Farrell and L. E. Jones, Isomorphism conjectures in algebraic K-theory, J. Amer. Math. Soc. 6 (1993), 249–297.

[Foss] R. Fossum, Vector bundles, Inv. Math. 8 (1969), 222–225..

[FS] E. Friedlander and A. Suslin, The spectral sequence relating algebraic K-theory to motivic cohomology, Ann. Sci. École Norm. Sup. 35 (2002), 773–875.

[Fuchs] L. Fuchs, Infinite Abelian Groups II, Academic Press, 1973.

[Fulton] W. Fulton, Intersection Theory, Ergebnisse Math., vol. 2, Springer-Verlag, 1984.

[FL] W. Fulton and S. Lang, Riemann-Roch Algebra, Springer-Verlag, 1985.

[Gabber] O. Gabber, K-theory of Henselian local rings and Henselian pairs, Letter to Karoubi, May 1983., Contemp. Math., vol. 126, 1992, pp. 59–70.

[Gabriel] P. Gabriel, Des catégories abeliennes,, Bull. Soc. Math. France 90 (1962), 323–448.

[GAGA] J.-P. Serre, Géométrie algébrique et géométrie analytique, Ann. Inst. Fourier 6 (1956), 1–42.

[GeiL] T. Geisser and M. Levine, The K-theory of fields in characteristic p, Invent. Math. 139 (2000), 459–493.

[GL01] T. Geisser and M. Levine, The Bloch-Kato conjecture and a theorem of Suslin-Voevodsky, J. Reine Angew. Math. 530 (2001), 55–103.

[Ger72] S. Gersten, Higher K-theory of rings, Lecture Notes in Math., vol. 341, Springer-Verlag, 1973, pp. 3–42.

[Ger74] S. Gersten, K-theory of free rings, Comm. Alg. 1 (1974), 39–64.

[GSz] P. Gille and T. Szamuely, Central Simple Algebras and Galois Cohomology, Cambridge Univ. Press, 2006.

[Gillet] H. Gillet, Riemann-Roch theorems for higher algebraic K-theory, Adv. in Math. 40 (1981), 203–289.

[GG] H. Gillet and D. Grayson, The loop space of the Q-construction, Ill. J. Math. 11 (1987), 574–597.

[GiL] H. Gillet and M. Levine, The relative form of Gersten’s conjecture over a discrete valuation ring: the smooth case, J. Pure Applied Algebra 46 (1987), 59–71.

[GiS] H. Gillet and C. Soulé, Filtrations on higher K-theory, Proc. Symp. Pure Math., vol. 67, 1999, pp. 89–148.

[Gdearl] K. Goodearl, Von Neumann regular rings, Pitman, 1979.

[Gdrl1] K. Goodearl, Torsion in K_{0} of unit-regular rings, Proc. Edinburgh Math. Soc. 38 (1995), 331-341.

[GR] H. Grauert and R. Remmert, Theory of Stein Spaces, Springer-Verlag, 1977.

[GQ] D. Grayson, Higher algebraic K-theory: II (after D. Quillen), Lecture Notes in Math., vol. 551, Springer-Verlag, 1976.

[GQ82] D. Grayson, Finite generation of K-groups of a curve over a finite field (after D. Quillen), Lecture Notes in Math., vol. 966, Springer-Verlag, 1982, pp. 69–90.

[Gra87] D. Grayson, Exact Sequences in algebraic K-theory, Illinois J. Math. 31 (1987), 598– 617.

[Gra88] D. Grayson, The K-theory of semilinear endomorphisms, J. Algebra 113 (1988), 358– 372.

[Gra95] D. Grayson, Weight filtrations via commuting automorphisms, K-Theory 9 (1995), 139–172.

[GiS] H. Gillet and C. Soulé, Filtrations on higher K-theory, Proc. Symp. Pure Math., vol. 67, 1999, pp. 89–148.

[GA] P. Griffiths and J. Adams, Topics in Algebraic and Analytic Geometry, Princeton U. Press, 1974.

[Groth57] A. Grothendieck, Sur la classification des fibrés holomorphes sur la sphere de Riemann, Amer. J. Math. 79 (1957), 121–138.

[GDix] A. Grothendieck, Classes de Chern et représentations linéaires des groupes discrets, Dix Exposés sur la Cohomologie des Schémas, North-Holland, 1968, pp. pp. 215.305.

[HW] C. Haesemeyer and C. Weibel, The Norm Residue Theorem in Motivic Cohomology, in preparation.

[HS] B. Harris and G. Segal, K_{i} groups of rings of algebraic integers, Annals of Math. 101 (1975), 20–33.

[Hart] R. Hartshorne, Algebraic Geometry, Springer-Verlag, 1977.

[H20] R. Hartshorne, Residues and Duality, Lecture Notes in Math., vol. 20, Springer-Verlag, 1966.

[HH] J.-C. Hausmann and D. Husemoller, Acyclic maps, Enseign. Math. 25 (1979), 53–75.

[HM] L. Hesselholt and I. Madsen, On the K-theory of local fields, Annals of Math. 158 (2003), 1-113.

[Hiller] H. Hiller, λ-rings and algebraic K-theory, J. Pure Applied Algebra 20 (1981), 241–266.

[Hovey] M. Hovey, Model categories, Math. Surveys Monographs 63, AMS, 1999.

[HSS] M. Hovey, B. Shipley and J. Smith, Symmetric spectra, J. AMS 13 (2000), 149–208.

[Hu] A. Huber, Realization of Voevodsky’s motives, J. Algebraic Geometry 9 (2000), 755- [Huse] D. Husemoller, Fibre Bundles, (Second Edition), Springer-Verlag, 1974.

[Igusa] K. Igusa, What happens to Hatcher and Wagoner’s formulas for π_{0}C(M) when the first Postnikov invariant of M is nontrivial?, Lecture Notes in Math., vol. 1046, Springer-Verlag, 1984, pp. 104–172.

[Izh] O. Izhboldin, On the torsion subgroup of Milnor’s K-groups, Soviet Math. Doklady 35 (1987), 493–495; On p-torsion in K_{∗}^{M} for fields of characteristic p., Algebraic K-theory, Adv. Soviet Math., vol. 4, Amer. Math. Soc., 1991, pp. 129–144.

[J37] N. Jacobson, p-alghebras of exponent p, Bull. AMS 43 (1937), 667–670.

[J] J. F. Jardine, Generalized étale cohomology theories, Birkhäuser, 1997.

[Kap58] I. Kaplansky, Projective modules, Annals of Math. 68 (1958), 372–377.

[Kar63] M. Karoubi, Les isomorphismes de Chern et de Thom-Gysin en K-théorie,, Exposé 16, Seminaire H. Cartan, 1963/64.

[KV71] M. Karoubi and O. Villamayor, K-theorie algebrique et K-theorie topologique, Math. Scand. 28 (1971), 265–307.

[Ka82] K. Kato, Symmetric bilinear forms, Quadratic forms and Milnor K-theory in characteristic two, Invent. Math. 66 (1982), 493–510.

[Ke90] B. Keller, Chain complexes and stable categories, Manuscripta Math. 67 (1990), 379– 417.

[Keu78] F. Keune, The relativization of K_{2}, J. Algebra 54 (1978), 159–177.

[Kum] E. Kummer, Collected Papers I (A. Weil, ed.), Springer Verlag, 1975.

[Kur] M. Kurihara, Some remarks on conjectures about cyclotomic fields and K-groups of ℤ, Compositio Math. 81 (1992), 223–236.

[Lam] T.Y. Lam, Serre’s Conjecture, Lecture Notes in Math., vol. 635, Springer-Verlag, 1978.

[Landsbg] S. Landsburg, Patching theorems for projective modules, J. Pure Appl. Alg. 21 (1981), 261–277.

[LSz] R. Lee and R. Szczarba, The group K_{3}(ℤ) is cyclic of order forty-eight, Annals Math. 104 (1976), 31–60.

[Le01] M. Levine, Techniques of localization in the theory of algebraic cycles, J. Alg. Geom. 10 (2001), 299–363.

[Le08] M. Levine, The homotopy coniveau tower, J. Topology 1 (2008), 217–267.

[Li] S. Lichtenbaum, On the values of zeta and L-functions: I, Annals of Math. 96 (1972), 338–360.

[Li2] , Values of zeta functions, étale cohomology, and algebraic K-theory, Lecture Notes in Math., vol. 342, Springer Verlag, 1973, pp. 489–501.

[Li3] , Values of zeta functions at non-negative integers, Lecture Notes in Math., vol. 1068, Springer Verlag, 1984, pp. 127–138.

[Lo76] J.-L. Loday, K-théorie algébrique et représentations de groupes, Annales Sci. ENS 9 (1976), 309–377.

[Lo78] J.-L. Loday, Cohomologie et groupe de Steinberg relatifs, J. Algebra54 (1978), 178–202.

[Mac] S. MacLane, Categories for the Working Mathematician, Springer-Verlag, 1971.

[Matsu] H. Matsumura, Commutative ring theory, Cambridge Univ. Press, 1986.

[May] J.P. May, Simplicial Objects in Algebraic Topology, Van Nostrand, 1967.

[May74] J.P. May, E_{∞} ring spaces, group completions and permutative categories, London Math. Soc. Lecture Notes, vol. 11, Cambridge Univ. Press, 1974, pp. 61–93.

[May80] J.P. May, Pairings of Categories and Spectra, J. Pure Applied Algebra (1980), 299-346.

[MTh] J.P. May and R. Thomason, The uniqueness of infinite loop space machines, Topology 17 (1978), 205–224.

[MVW] C. Mazza, V. Voevodsky and C. Weibel, Lecture Notes in Motivic Cohomology, Clay Math. Monographs 4, AMS, 2006.

[MM82] P. Menal and J. Moncasi, On regular rings with stable range 2, J. Pure Applied Algebra 22 (1982), 25–40.

[Merk] A. Merkurjev, On the torsion in K_{2} of local fields, Annals of Math. 118 (1983), 375– 381.

[MS] A. Merkurjev and A. Suslin, K-cohomology of Severi-Brauer Varieties and the Norm Residue Homomorphism, Math. USSR Izvestiya 21 (1983), 307–340.

[MS2] A. Merkurjev and A. Suslin, On the norm residue homomorphism of degree three, Izv. Akad. Nauk SSSR Ser. Mat. 54 (1990), 339–356.

[Milne] J. Milne, Étale Cohomology, Princeton Univ. Press, 1980.

[Milne2] J. Milne, Arithmetic Duality Theorems, Academic Press, 1986. [M-QF] J. Milnor, Algebraic K-theory and Quadratic Forms, Invent. Math. 9 (1970), 318–344.

[Milnor] J. Milnor, Introduction to Algebraic K-theory, Annals of Math. Study 72, Princeton Univ. Press, 1971.

[M-Lie] J. Milnor, On the homology of Lie groups made discrete, Comment. Math. Helv. 58

(1983), 72–85.

[M-SBF] J. Milnor and D. Husemoller, Symmetric Bilinear Forms, Springer-Verlag, 1973.

[MM] J. Milnor and J. Moore, On the structure of Hopf algebras, Annals of Math. 81 (1965), 211–264.

[MSt] J. Milnor and J. Stasheff, Characteristic classes, Annals of Math. Study, vol. 76, Princeton Univ. Press, 1974.

[MimTo] M. Mimura and H. Toda, Topology of Lie Groups, Translations of Math. Monographs, vol. 91, AMS, 1991.

[Mit] S. Mitchell, Hypercohomology spectra and Thomason’s descent theorem, Fields Inst. Commun. vol. 16, Algebraic K-theory (Toronto, ON, 1996), Amer. Math. Soc., Providence, RI, 1997, pp. 221–277.

[MV] F. Morel and V. Voevodsky, 𝔸^{1}-homotopy theory of schemes, Publ. IHES 90 (2001), 45–143.

[Mum] D. Mumford, Lectures on curves on an algebraic surface, Annals Studies in Math., vol. 59, Princeton University Press, Princeton, 1966.

[MmAB] D. Mumford, Abelian Varieties, Oxford U. Press, 1970.

[Nak] M. Nakaoka, Decomposition Theorem for Homology Groups of Symmetric Groups, Annals of Math. 71 (1960), 16–42..

[N] J. Neisendorfer, Primary homotopy theory, Memoirs AMS, vol. 232, 1980.

[Nen] A. Nenashev, K_{1} by generators and relations, J. Pure Appl. Alg. 131 (1998), 195–212..

[NS] Y. Nesterenko and A. Suslin, Homology of the general linear group over a local ring, and Milnor’s K-theory, Izv. Akad. Nauk SSSR Ser. Mat. 53 (1989), 121–146.

[Neu] J. Neukirch, Algebraic number theory, Translated from the 1992 German original, Springer-Verlag, 1999.

[Oliver] R. Oliver, Whitehead Groups of Finite Groups, Cambridge Univ. Press, 1988.

[OMeara] T. O’Meara, Introduction to quadratic forms, reprinted in 1971 and 2000, Springer, 1963.

[OVV] D. Orlov, A. Vishik and V. Voevodsky, An exact sequence for K_{∗}^{M} /2 with applications to quadratic forms, Annals Math. 165 (2007), 1–13.

[Pa] I. Panin, On a theorem of Hurewicz and K-theory of complete discrete valuation rings, Math. USSR Izvestiya 29 (1987), 119–131.

[Puppe] D. Puppe, A theorem on semi-simplicial monoid complexes, Annals Math. 70 (1959), 379–394.

[Q66] D. Quillen, Spectral sequences of a double semi-simplicial group, Topology 5 (1966), 155–157.

[Q70] D. Quillen, On the cohomology and K-theory of the general linear groups over a finite field, Revised version of 1970 preprint “the K-theory associated to a finite field”, Annals of Math. 96 (1972), 552–586.

[Q341] D. Quillen, Higher algebraic K-theory: I, Lecture Notes in Math., vol. 341, Springer-Verlag, 1973, pp. 85–147.

[Q73] D. Quillen, Finite generation of the groups K_{i} of rings of algebraic integers, Lecture Notes in Math., vol. 341, Springer-Verlag, 1973, pp. 195–214.

[Qlet] D. Quillen, Letter from Quillen to Milnor on ℑ(π_{i}O → π_{i}^{S} → K_{i}ℤ), Lecture Notes in Math., vol. 551, Springer Verlag, 1976, pp. 182–188.

[Q76] D. Quillen, Characteristic classes of representations, Lecture Notes in Math., vol. 551, Springer-Verlag, 1976, pp. 189–216.

[Reid] L. Reid, N-dimensional rings with an isolated singular point having nonzero K_{−N}, K-Theory 1 (1987), 197–205.

[Rognes] J. Rognes, K_{4}(ℤ) is the trivial group, Topology 39 (2000), 267–281.

[ROst] J. Rognes and P. A. Østvaer, Two-primary algebraic K-theory of two-regular number fields, Math. Zeit. 233 (2000), 251–263.

[RW] J. Rognes and C. Weibel, Two-primary algebraic K-theory of rings of integers in number fields, Jour. AMS 13 (2000), 1–54.

[Ro96] J. Rosenberg, The algebraic K-theory of operator algebras, K-theory 12 (1997), 75–99.

[Schl] M. Schlichting, Delooping the K-theory of exact categories, Topology 43 (2004), 189– 1103.

[Segal] G. Segal, Categories and cohomology theories, Topology 13 yr1974, 293–312. [S-FAC] J.-P. Serre, Faisceaux algébriques cohérents, Annals of Math. 61 (1955), 197–278. [S-LF] J.-P. Serre, Local Fields, Springer-Verlag, 1979.

[SGA4] M. Artin, A. Grothendieck and J.-L. Verdier,, Théorie des topos et cohomologie étale des schemas, Lecture Notes in Math., vol. 269, 270 and 305, Springer, 1972–73.

[SGA6] P. Berthelot, A. Grothendieck and L. Illusie, Théorie des intersections et théorème de Riemann-Roch, Lecture Notes in Math., vol. 225, Springer, 1971.

[Shatz] S. Shatz, Profinite grups, arithmetic, and geometry, Annals of Math. Study, vol. 67, Princeton Univ. Press, 1972.

[Siv] I. Sivitskii, Torsion in Milnor K-groups for a local field, Mat. Sb. 126 (168) (1985), 576–583..

[Sou] C. Soulé, K-théorie des anneaux d’entiers de corps de nombres et cohomologie étale, Invent. Math. 55 (1979), 251–295.

[So78] , Addendum to the article “On the torsion in K_{∗}(ℤ)” by R. Lee and R. Szczarba, Duke Math. J. 45 (1978), 131–132.

[Sou81] , On higher p-adic regulators, Lecture Notes in Math., vol. 854, Springer Verlag, 1981, pp. 372–401.

[Sou85] C. Soulé, Opérations en K-théorie algébrique,, Canad. J. Math. 37 (1985), 488–550.

[Staf] R. Staffeldt, On Fundamental Theorems of Algebraic K-theory, K-theory 1 (1989), 511–532.

[St82] J. Stienstra, Operations in the higher K-theory of endomorphisms, CMS Conf. Series, vol. 2, AMS, 1982, pp. 59–115.

[St85] J. Stienstra, Correction to Cartier-Dieudonné theory for Chow groups, J. Reine Angew. Math. 365 (1985), AMS, 218–220.

[Su76] A. Suslin, On a theorem of Cohn, Zap. Nauk LOMI (1976), 127–130; English transl. in J. Soviet Math. 17 (1981), 1801–1803.

[Su77] A. Suslin, On the structure of the special linear group over polynomial rings, Math. USSR Izv. 11 (1977), 221–238.

[Su81] A. Suslin, On the equivalence of K-theories, Comm. in Alg. 9 (1981), 1559–1565.

[S82] A. Suslin, Stability in algebraic K-theory, Lecture Notes in Math., vol. 966, Springer-Verlag, 1982, pp. 304–333.

[Su83] A. Suslin, On the K-theory of algebraically closed fields, Invent. Math. 73 (1983), 241–245.

[Su-KM] A. Suslin, Homology of GL_{n}, characteristic classes and Milnor K-theory, Lecture

Notes in Math., vol. 1046, Springer Verlag, 1984, pp. 357–375.

[Su84] A. Suslin, On the K-theory of local fields, J. Pure Appl. Alg. 34 (1984), 319–330.

[Su86] A. Suslin, Algebraic K-theory of fields, Proc. Berkeley ICM Vol. 1, Amer. Math. Soc., Providence, 1987, pp. 222–244.

[Su87] A. Suslin, Torsion in K_{2} of fields, K-theory 1 (1987), 5–29.

[Su91] A. Suslin, K_{3} of a field, and the Bloch group, Trudy Mat. Inst. Steklov 183 (1990), 180–199 (Russian); English transl. in Proc. Steklov Inst. Math. 1991, 217–239.

[Su95] A. Suslin, Excision in integer algebraic K-theory, Trudy Mat. Inst. Steklov 208 (1995), 290–317 (Russian); English transl. in Teor. Chisel, Algebra i Algebr. Geom..

[Su03] A. Suslin, On the Grayson spectral sequence, Trudy Mat. Inst. Steklova 241 (2003), 218–253; English transl. in Proc. Steklov Inst. Math. 241 (2003), 202–237.

[SV00] A. Suslin and V. Voevodsky, Bloch-Kato conjecture and motivic cohomology with finite coefficients, NATO Sci. Ser. C Math. Phys. Sci., vol. 548, Kluwer, 2000, pp. 117–189.

[SuW] A. Suslin and M. Wodzicki, Excision in algebraic K-theory, Annals of Math. 136 (1992), 51–122.

[Suz] M. Suzuki, Group Theory I, Springer-Verlag, 1982.

[Swan] R. Swan, Algebraic K-theory, Lecture Notes in Math., vol. 76, Springer, 1968.

[Swan70] R. Swan, A Splitting Principle in Algebraic K-theory, Proc. Sympos. Pure Math., vol. XXI, AMS, 1971, pp. 155–159.

[Swan71] R. Swan, Excision in algebraic K-theory, J. Pure Applied Algebra 1 (1971), 221–252..

[Swan72] R. Swan, Some relations between higher K-functors, J. Algebra 21 (1972), 113–136.

[Swan80] R. Swan, On Seminormality, J. Algebra 67 (1980), 210–229.

[Tate] J. Tate, On the torsion in K_{2} of fields, Algebraic Number Theory (1976 Kyoto Internat. Sympos.), Japan Soc. Promotion Sci., Tokyo, 1977.

[TT] R.W. Thomason and T. Trobaugh, Higher algebraic K-theory of schemes and of derived categories, The Grothendieck Festschrift III, Progress in Math., vol. 88, Birkhäuser, 1990, pp. 247–435.

[Totaro] B. Tot, Milnor K-theory is the simplest part of algebraic K-theory, K-Theory 6 (1992), 177–189.

[Van29] H. Vandiver, On Fermat’s Last Theorem, Trans. AMS 31 (1929), 613–642.

[Van34] Fermat’s Last Theorem and the second factor in the cyclotomic class number, Bull. AMS 40 (1934), 118–126.

[Van46] H. Vandiver, Fermat’s Last Theorem: its history and the nature of the known results concerning it, Amer. Math. Monthly 53 (1946), 555–578.

[V69] L. Vaserstein, On the stabilization of the general linear group over a ring, Math USSR Sb. 8 (1969), 383–400.

[V76] L. Vaserstein, Foundations of algebraic K-theory, Uspehi Mat. Nauk 31:4 (1976), 87– 149 (Russian); English transl. in Russian Math. Surveys 31:4 (1976), 89–156.

[V04] L. Vaserstein, On the Whitehead determinant for semi-local rings, J. Algebra 283 (2005), 690–699.

[Verd] J.-L. Verdier, Catégories dérivées, Cohomologie étale (SGA4-1/2), Lecture Notes in Math., vol. 569, Springer-Verlag, 1977.

[VV02a] V. Voevodsky, Open problems in the motivic stable homotopy theory I, Motives, polylogarithms and Hodge Theory I, 2002, pp. 3–34.

[VV02b] V. Voevodsky, A possible new approach to the motivic spectral sequence for algebraic K-theory, Contemp. Math., vol. 293, 2002, pp. 371–379.

[V-MC] V. Voevodsky, Motivic cohomology with ℤ/2-coefficients, Publ. Math. IHES 98 (2003),

59–104.

[Vo] A. Vorst, Localization of the K-theory of Polynomial Extensions, Math. Ann. 244 (1979), 33–53.

[Wag] J. B. Wagoner, Continuous cohomology and p-adic K-theory, Lecture Notes in Math., vol. 551, Springer Verlag, 1976, pp. 241–248.

[Wa78] F. Waldhausen, Algebraic K-theory of generalized free products, Annals of Math. 108 (1978), 135–256.

[W1126] F. Waldhausen, Algebraic K-theory of spaces, Lecture Notes in Math., vol. 1126, Springer-Verlag, 1985.

[Wash] L. Washington, Introduction to Cyclotomic Fields, Springer, 1982.

[We80] C. Weibel, K-theory and Analytic Isomorphisms, Invent. Math 61 (1980), 177–197.

[We81] C. Weibel, A survey of products in algebraic K-theory, Lecture Notes in Math., vol. 854, Springer-Verlag, 1981, pp. 494–517.

[We-Az] C. Weibel, K-theory of Azumaya algebras, Proc. AMS 81 (1981), 1–7.

[We89] C. Weibel, Homotopy algebraic K-theory, Contemporary Math., vol. 83, AMS, 1989, pp. 461–488.

[We91] C. Weibel, Pic is a contracted functor, Invent. Math. 103 (1991), 351–377.

[We93] C. Weibel, Étale Chern classes at the prime 2, NATO Sci. Ser. C Math. Phys. Sci., vol. 407, Kluwer, 1993, pp. 249–286.

[WHomo] C. Weibel, An introduction to homological algebra, Cambridge Univ. Press, 1994.

[We06] C. Weibel, Higher wild kernels and divisibility in the K-theory of number fields, J. Pure Appl. Alg. 206 (2006), 222–244.

[Wh50] J.H.C. Whitehead, Simple Homotopy Types, Amer. J. Math 72 (1950), 1–57.

[Wh] G.W. Whitehead, Elements of Homotopy Theory, Springer-Verlag, 1978.

[Wi] A. Wiles, The Iwasawa conjecture for totally real fields, Annals of Math. 131 (1990), 493–540.